# The stability of propagating plane inertial waves in rotating fluids

### v1

@@ Acknowledgements.
AA was supported by a Leverhulme Trust Early Career Fellowship and by STFC grant ST/S000275/1. AJB was supported by STFC grants ST/S000275/1, ST/W000873/1 and UKRI1179. We would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programmes DYT2 and “Anti-diffusive dynamics: from sub-cellular to astrophysical scales”, supported by EPSRC grant no EP/R014604/1. We would like to thank the Kavli Summer Program In Astrophysics 2021: Fluid dynamics of the Sun and Stars, during which this project was initiated.
@@
Appendix A Energy conservation equation for Floquet system
The evolution equation for the kinetic energy of the perturbations $K^{\prime}(\phi)=\left(|\hat{u}|^{2}+|\hat{v}|^{2}+|\hat{w}|^{2}\right)/2$ is obtained by first using the Floquet ansatz (Eq. (14)) in the momentum equation (Eqs. 10, 11, and 12) to express it in terms of periodic $\hat{f}$ variables, and then multiplying by the complex conjugates of the velocity components $\hat{u}^{*}$, $\hat{v}^{*}$, and $\hat{w}^{*}$, respectively. This gives
$2\mathrm{Re}[\sigma]K^{\prime}+W^{\prime}\frac{\mathrm{d}K^{\prime}}{\mathrm{d}\phi}=$ $-\mathrm{Re}[\hat{u}^{*}\hat{w}]\frac{\mathrm{d}U^{\prime}}{\mathrm{d}\phi}-\mathrm{Re}[\hat{v}^{*}\hat{w}]\frac{\mathrm{d}V^{\prime}}{\mathrm{d}\phi}-\mathrm{Im}\left[\alpha\hat{p}^{*}\hat{u}+\beta\hat{p}^{*}\hat{v}+\gamma\hat{p}^{*}\hat{w}\right]-\mathrm{Re}\left[\frac{\mathrm{d}\hat{p}^{*}}{\mathrm{d}\phi}\hat{w}\right$]
@@ Appendix A Energy conservation equation for Floquet system
where $\mathrm{Re}$ and $\mathrm{Im}$ are the real and imaginary parts. This equation is similar to the energy transfer equation for propagating IGWs derived in Lombard and Riley [32, Eq. (20)], except without buoyancy fluxes. The first two terms on the right-hand side of Eq. (19) are shear production or extraction terms, which can act to transfer energy between the primary wave and the perturbations. By contrast, the advective term on the left-hand side ($\propto W^{\prime}$) and the two pressure terms on the right-hand side are conservative and vanish when integrated over one wavelength. Carrying out the integration gives the relation between the growth rate and the various source and sink terms:
$2\mathrm{Re}[\sigma]=-\left\langle\mathrm{Re}[\hat{u}^{*}\hat{w}]\frac{\mathrm{d}U^{\prime}}{\mathrm{d}\phi}\right\rangle_{\phi}-\left\langle\mathrm{Re}[\hat{v}^{*}\hat{w}]\frac{\mathrm{d}V^{\prime}}{\mathrm{d}\phi}\right\rangle_{\phi}+\mathrm{Ek}\,\left\langle\mathrm{Re}\left[\hat{u}^{*}L_{2}\hat{u}+\hat{v}^{*}L_{2}\hat{v}+\hat{w}^{*}L_{2}\hat{w}\right]\right\rangle_{\phi},$ (20)
where we have normalized by the average kinetic energy density $\left\langle K^{\prime}\right\rangle_{\phi}$ with $\langle\cdot\rangle_{\phi}=\int^{2\pi}_{0}\cdot\,\mathrm{d}\phi\,/(2\pi\left\langle K^{\prime}\right\rangle_{\phi})$. The growth rate can be positive when the shear terms are positive (extracting energy from the primary wave) and the viscous dissipation term is relatively small.
@@ Appendix B Theory for small amplitude waves: TRI and PSI
$\bm{u}_{0}=\sum_{X}X(\tau)\bm{u}_{X}E_{X}(\bm{x},t)+c.c.,\quad E_{X}\equiv\mathrm{e}^{i(-\omega_{X}t+\bm{k}_{X}\cdot\bm{x})},$ (29)
where $c.c.$ denotes the complex conjugate and $X$ denotes a given wave. Each wave satisfies the dispersion relation
@@
Figure 10: The fastest growing unstable perturbations according to the PSI and the Floquet analysis for various amplitudes $A^{\prime}$ of a primary IW with fixed frequency $\omega/2\Omega=0.71$ ($\theta=45^{\circ}$). Panel (a) shows $\sigma^{\max}$ versus $\alpha$ at fixed $\beta=0$ and $\gamma=0$. Panel (b) shows $\sigma^{\max}$ versus $\beta$ at fixed $\alpha=0$ and $\gamma=0$. The PSI and Floquet theories agree at low wave amplitudes. However, they deviate at larger amplitudes when couplings of the $n$th mode with modes beyond $n\pm 1$ become significant in the Floquet theory, which invalidates the the three-wave assumption made in the PSI analysis. The Floquet solutions for $A^{\prime}=0.3$ in panels (a) and (b) are identical to a horizontal slice at $\beta=0$ and a vertical slice at $\alpha=0$, respectively, through panel (b) in Figure 1. The fastest growth rate of the PSI, $\sigma_{\rm PSI}\approx 0.325,$ is shown as the dashed, gray line, which closely matches the growth rates of the perturbations with $\alpha=0$ in panel (b) (see Section II.2 for discussion).
@@
$+(\partial_{\tau}b)\bm{u}_{b}E_{b}+\mathrm{i}b\left[(\bm{\hat{U}}_{p}\cdot\bm{k}_{b})\bm{u}_{b}E_{p}+(\hat{\bm{U}}^{*}_{p}\cdot\bm{k}_{b})\bm{u}_{b}E_{p}^{*}+(\bm{u}_{b}\cdot\bm{k}_{p})\hat{\bm{U}}_{p}E_{p}-(\bm{u}_{b}\cdot\bm{k}_{p})\hat{\bm{U}}^{*}_{p}E_{p}^{*}\right]E_{b}+c.c.$ (34)
@@
The solvability condition requires the components of $\bm{g}_{a}$ and $\bm{g}_{b}$ to vanish, giving
@@
where $C_{ab}$ and $C_{ba}$ are determined from the coefficients in $\bm{g}_{a}$ and $\bm{g}_{b}$ (i.e., Equation (36)). An exponential solution $a,b\propto\exp(\sigma\tau)$ has growth rate [^1] In the cases discussed below and in Fig. 10, where either $k_{a,x}=\alpha=0$ or $k_{a,y}=\beta=0$, we found that the coefficients $C_{ab}$ and $C_{ba}$ (and hence $\sigma$) are both real. We have also confirmed this numerically for selected values of $\alpha,\beta$ and $\gamma$ when evaluating the growth rate based on the analysis in this appendix. $\sigma=\pm\sqrt{C_{ab}C_{ba}}$ such that the dimensional growth rate is $\sigma=\pm\sqrt{C_{ab}C_{ba}}A^{\prime}|\omega|$, where we have equated $\epsilon$ with the primary wave amplitude. The derivation leading to Eq. 37 is straightforwardly valid as long as the coefficient of $\partial_{\tau}a$ in Eq. (36) does not vanish, in which case we must consider it more carefully. This coefficient reduces to $2\mathrm{i}\tilde{\omega}_{a}k_{a}^{2}\bm{u}_{a}$, which vanishes when $\omega_{a}=k_{a,\varphi}W$, corresponding to the perturbation comoving with the primary wave, somewhat like a “corotation resonance”.
@@
For small $A^{\prime}$, we expect that the fastest growing instabilities will be those for which the resonance conditions in Eq. (33) are most precisely satisfied, which is generally easier at large $\beta^{2}$ and $n^{2}$. In this limit with $\bm{k}_{a}\approx\bm{k}_{b}$, the TRI reduces to the PSI where the daughter waves have half the frequency of the primary wave in the rotating frame. For simplicity, and motivated by our numerical results, we consider modes with $\alpha=0$ and $s=1=k_{z}/|k_{z}|$ (so W=1W=-1) such that $\bm{k}_{a}=(0,\beta,n)$ and $\bm{k}_{b}=(0,\beta,n+1)$, thus satisfying the resonance condition Eq. (33), with $k_{a}^{2}=\beta^{2}+n^{2}$ and $k_{b}^{2}=\beta^{2}+(n+1)^{2}$. The eigenvectors ($X=a,b$) simplify to
@@
$\bm{u}_{X}=\left(1,-\frac{\mathrm{i}\tilde{\omega}_{X}k_{X,\phi}}{\bm{\Omega}^{\prime}\cdot\bm{k}_{X}},\frac{\mathrm{i}\tilde{\omega}_{X}k_{X,y}}{\bm{\Omega}^{\prime}\cdot\bm{k}_{X}}\right)=\left(1,-\mathrm{i}\tilde{\omega}_{X},\pm\frac{ik_{X,y}}{k_{X}}\right),$ (38)
@@
From the dispersion relation Eq. (30), the frequencies are given by $\tilde{\omega}_{a}^{2}=n^{2}/k_{a}^{2}$ and $\tilde{\omega}_{b}^{2}=(n+1)^{2}/k_{b}^{2}$ since $\bm{\Omega}^{\prime}\cdot\bm{k}_{a}=n$ and $\bm{\Omega}^{\prime}\cdot\bm{k}_{b}=n+1$. To satisfy the resonance conditions, we decide to set $\tilde{\omega}_{a}=-n/k_{a}$ and $\tilde{\omega}_{b}=(n+1)/k_{b}$. The expressions for $C_{ab}$ and $C_{ba}$, and hence $\sigma$, can then be obtained as a function of $\beta$ and $n$. The large wavenumber limit of the product $C_{ab}C_{ba}$ can be taken by fixing the wavevector orientation $r=\beta/n$ for a constant ratio $r$ and then taking the limit $\beta\to\infty$ (or equivalently, $n\to\infty$). We obtain
$\sqrt{C_{ab}C_{ba}}\to\frac{1}{4}\left(\frac{r}{1+r^{2}}+\frac{r}{\sqrt{1+r^{2}}}\right),$ (39)
@@
This growth rate for PSI is consistent with the results obtained in [39, 40]. They find $\sigma_{\mathrm{PSI}}=0.25\sin\alpha_{2}(1+\cos\alpha_{2})A^{\prime}|\omega|=(3\sqrt{3}/16)A^{\prime}|\omega|$ with the inclination angle of the wavevector $\alpha_{2}=\pi/3$ corresponding to $\tan{\alpha_{2}}=r=\sqrt{3}$. Hence, our results are in agreement. Note that this PSI limit does not necessarily lead to the fastest growth rate over all possible modes – in that we have not proven that it is the optimum growth rate over all $\beta$ and $n$, though we expect this to be the case.

### v2

@@ Acknowledgements.
AA was supported by a Leverhulme Trust Early Career Fellowship and by STFC grant ST/S000275/1. AJB was supported by STFC Grant No. ST/S000275/1, ST/W000873/1 and UKRI1179. We would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programs DYT2 and “Anti-diffusive dynamics: from sub-cellular to astrophysical scales”, supported by EPSRC Grant No. EP/R014604/1. We would also like to thank the Kavli Foundation and the Max Plank Institute for Solar System Research at the University of Goettingen who supported the “Kavli Summer Program in Astrophysics 2021: Fluid Dynamics of the Sun and Stars,” during which this project was initiated.
@@
Appendix A Energy conservation equation for the Floquet system
The kinetic energy evolution of the perturbations $K^{\prime}(\phi)=\left(|\hat{u}|^{2}+|\hat{v}|^{2}+|\hat{w}|^{2}\right)/2$ is obtained by using the Floquet ansatz (Eq. (14)) in the momentum equation (Eqs. 10, 11, and 12) to express it in terms of periodic $\hat{f}$ variables, and then multiplying by the complex conjugates of the velocity components $\hat{u}^{*}$, $\hat{v}^{*}$, and $\hat{w}^{*}$, respectively. This gives
$2\mathrm{Re}[\sigma^{\prime}]K^{\prime}+W^{\prime}\frac{\mathrm{d}K^{\prime}}{\mathrm{d}\phi}=$ $-\mathrm{Re}[\hat{u}^{*}\hat{w}]\frac{\mathrm{d}U^{\prime}}{\mathrm{d}\phi}-\mathrm{Re}[\hat{v}^{*}\hat{w}]\frac{\mathrm{d}V^{\prime}}{\mathrm{d}\phi}-\mathrm{Im}\left[\alpha\hat{p}^{*}\hat{u}+\beta\hat{p}^{*}\hat{v}+\gamma\hat{p}^{*}\hat{w}\right]-\mathrm{Re}\left[\frac{\mathrm{d}\hat{p}^{*}}{\mathrm{d}\phi}\hat{w}\right$]
@@ Appendix A Energy conservation equation for the Floquet system
This equation is similar to the energy transfer equation for propagating IGWs derived in Lombard and Riley [47, Eq. (20)], except without buoyancy fluxes. The first two terms on the right-hand side of Eq. (19) are shear production or extraction terms, which can act to transfer energy between the primary wave and the perturbations. By contrast, the advective term on the left-hand side ($\propto W^{\prime}$) and the two pressure terms on the right-hand side are conservative and vanish when integrated over one wavelength. Carrying out the integration gives the relation between the growth rate and the various source and sink terms:
$2\mathrm{Re}[\sigma^{\prime}]=-\left\langle\mathrm{Re}[\hat{u}^{*}\hat{w}]\frac{\mathrm{d}U^{\prime}}{\mathrm{d}\phi}\right\rangle_{\phi}-\left\langle\mathrm{Re}[\hat{v}^{*}\hat{w}]\frac{\mathrm{d}V^{\prime}}{\mathrm{d}\phi}\right\rangle_{\phi}+\mathrm{Ek}\,\left\langle\mathrm{Re}\left[\hat{u}^{*}L_{2}\hat{u}+\hat{v}^{*}L_{2}\hat{v}+\hat{w}^{*}L_{2}\hat{w}\right]\right\rangle_{\phi},$ (20)
where $\langle\cdot\rangle_{\phi}=\int^{2\pi}_{0}\cdot\,\mathrm{d}\phi\,/(2\pi\left\langle K^{\prime}\right\rangle_{\phi})$. The growth rate can be positive when the shear terms are positive (extracting energy from the primary wave), and the viscous dissipation term is relatively small.
@@ Appendix B Theory for small amplitude waves: TRI and PSI
$\bm{u}_{0}=\sum_{X}X(\tau)\bm{u}_{X}E_{X}(\bm{x},t)+\mathrm{c.c.},\quad E_{X}\equiv\mathrm{e}^{\mathrm{i}(-\omega_{X}t+\bm{k}_{X}\cdot\bm{x})},$ (29)
where $\mathrm{c.c.}$ denotes the complex conjugate and $X$ denotes a given wave. Each wave satisfies the dispersion relation
@@
Figure 11: The fastest growing unstable perturbations according to the PSI and the Floquet analysis for various amplitudes $A^{\prime}$ of a primary IW with fixed frequency $\omega/2\Omega=0.71$ ($\theta=45^{\circ}$). Panel (a) shows $\sigma^{\max}$ versus $\alpha$ at fixed $\beta=0$ and $\gamma=0$. Panel (b) shows $\sigma^{\max}$ versus $\beta$ at fixed $\alpha=0$ and $\gamma=0$. The PSI and Floquet theories agree at low wave amplitudes. However, they deviate at larger amplitudes when couplings of the $n\textendashth$ mode with modes beyond $n\pm 1$ become significant in the Floquet theory, which invalidates the three-wave assumption made in the PSI analysis. The Floquet solutions for $A^{\prime}=0.3$ in panels (a) and (b) are identical to a horizontal slice at $\beta=0$ and a vertical slice at $\alpha=0$, respectively, through panel (b) in Figure 1. The fastest growth rate of the PSI, $\sigma_{\rm PSI}\approx 0.325,$ is shown as the dashed, gray line, which closely matches the growth rates of the perturbations with $\alpha=0$ in panel (b) (see Section II.2 for discussion).
@@
$+(\partial_{\tau}b)\bm{u}_{b}E_{b}+\mathrm{i}b\left[(\bm{\hat{U}}_{p}\cdot\bm{k}_{b})\bm{u}_{b}E_{p}+(\hat{\bm{U}}^{*}_{p}\cdot\bm{k}_{b})\bm{u}_{b}E_{p}^{*}+(\bm{u}_{b}\cdot\bm{k}_{p})\hat{\bm{U}}_{p}E_{p}-(\bm{u}_{b}\cdot\bm{k}_{p})\hat{\bm{U}}^{*}_{p}E_{p}^{*}\right]E_{b}+\mathrm{c.c.},$ (34)
@@
The solvability condition requires each of the Cartesian components of the vectors $\bm{g}_{a}$ and $\bm{g}_{b}$ to vanish (so that $\bm{G}(\bm{u}_{0})$ contains no terms proportional to $E_{a}$ or $E_{b}$), giving
@@
where $C_{ab}$ and $C_{ba}$ are determined from the coefficients in $\bm{g}_{a}$ and $\bm{g}_{b}$ (i.e., Equation (36)). An exponential solution $a,b\propto\exp(\sigma^{\prime}\tau)$ has growth rate  [^1] In the cases discussed below and in Fig. 11, where either $k_{a,x}=\alpha=0$ or $k_{a,y}=\beta=0$, we found that the coefficients $C_{ab}$ and $C_{ba}$ (and hence $\sigma$) are both real. We have also confirmed this numerically for selected values of $\alpha,\beta$ and $\gamma$ when evaluating the growth rate based on the analysis in this appendix. $\sigma^{\prime}=\pm\sqrt{C_{ab}C_{ba}}$ such that the dimensional growth rate is $\sigma=\pm\sqrt{C_{ab}C_{ba}}A^{\prime}|\omega|$, where we have equated $\epsilon$ with the primary wave amplitude. The derivation leading to Eq. 37 is straightforwardly valid as long as the coefficient of $\partial_{\tau}a$ in Eq. (36) does not vanish, in which case it must be considered more carefully. This coefficient reduces to $2\mathrm{i}\tilde{\omega}_{a}k_{a}^{2}\bm{u}_{a}$, which vanishes when $\omega_{a}=k_{a,\varphi}W$, corresponding to the perturbation comoving with the primary wave, somewhat like a “critical layer” or “corotation resonance”.
@@
We expect that for small $A^{\prime}$, the maximum growth rate is achieved for modes that most precisely satisfy the resonance conditions in Eq. (33), which is generally easier at large perturbation wavenumbers $\alpha^{2}+\beta^{2}+(\gamma+n)^{2}\gg 1$. Here, we connect with the notation in the Floquet analysis by using $\alpha,\beta$ for the horizontal components of the daughter wave’s wave vector and $\gamma+n$ for the component along $\hat{\phi}$ (where $n$ is an integer and $0\leq\gamma<1$). In this limit with $\bm{k}_{a}\approx\bm{k}_{b}$, the TRI reduces to the PSI where the daughter waves have half the frequency of the primary wave in the rotating frame. For simplicity, and motivated by the numerical results, we consider modes with $\alpha=0$ and $s=1=k_{z}/|k_{z}|$ (so W=1W=-1) such that $\bm{k}_{a}=(0,\beta,\gamma+n)$ and $\bm{k}_{b}=(0,\beta,\gamma+n+1)$, thus satisfying the resonance condition Eq. (33), with $k_{a}^{2}=\beta^{2}+(\gamma+n)^{2}$ and $k_{b}^{2}=\beta^{2}+(\gamma+n+1)^{2}$. The eigenvectors ($X=a,b$) simplify to
@@
$\bm{u}_{X}=\left(1,-\frac{\mathrm{i}\tilde{\omega}_{X}k_{X,\phi}}{\bm{\Omega}^{\prime}\cdot\bm{k}_{X}},\frac{\mathrm{i}\tilde{\omega}_{X}k_{X,y}}{\bm{\Omega}^{\prime}\cdot\bm{k}_{X}}\right)=\left(1,-\mathrm{i}\tilde{\omega}_{X},\pm\frac{ik_{X,y}}{k_{X}}\right).$ (38)
@@
From the dispersion relation Eq. (30), the frequencies are given by $\tilde{\omega}_{a}^{2}=(\gamma+n)^{2}/k_{a}^{2}$ and $\tilde{\omega}_{b}^{2}=(\gamma+n+1)^{2}/k_{b}^{2}$ since $\bm{\Omega}^{\prime}\cdot\bm{k}_{a}=\gamma+n$ and $\bm{\Omega}^{\prime}\cdot\bm{k}_{b}=\gamma+n+1$. To satisfy the resonance conditions, we decide to set $\tilde{\omega}_{a}=-(\gamma+n)/k_{a}$ and $\tilde{\omega}_{b}=(\gamma+n+1)/k_{b}$. The expressions for $C_{ab}$ and $C_{ba}$, and hence $\sigma^{\prime}$, can then be obtained as a function of $\beta$ and $\gamma+n$. The large wavenumber limit of the product $C_{ab}C_{ba}$ is taken by fixing the wave vector orientation $r=\beta/(\gamma+n)$ for a constant ratio $r$ and then taking the limit $\beta\to\infty$ (or equivalently, $n\to\infty$). We obtain
$\sqrt{C_{ab}C_{ba}}\to\frac{1}{4}\left(\frac{r}{1+r^{2}}+\frac{r}{\sqrt{1+r^{2}}}\right).$ (39)
@@
This growth rate for PSI is consistent with the results obtained in [52, 1]. They find $\sigma_{\mathrm{PSI}}=0.25\sin\alpha_{2}(1+\cos\alpha_{2})A^{\prime}|\omega|=(3\sqrt{3}/16)A^{\prime}|\omega|$ with the inclination angle of the wavevector $\alpha_{2}=\pi/3$ corresponding to $\tan{\alpha_{2}}=r=\sqrt{3}$. Hence, our results are in agreement. This multiple-scale analysis is also numerically consistent with the Floquet theory as shown in Figure 11. Note that this PSI limit does not necessarily lead to the fastest growth rate over all possible modes – in that we have not proven that it is the optimum growth rate over all $\beta$ and $n$, though we expect this to be the case on physical grounds.